# Transformers for Stratified Spectropolarimetric Inversion: Proof of Concept

### v1

@@ Abstract
Solar spectropolarimetric inversion—inferring atmospheric conditions from the Stokes vector—is a key diagnostic tool for understanding solar magnetism, but traditional inversion methods are computationally expensive and sensitive to local minima. Advances in artificial intelligence (AI) offer faster solutions, but are often restricted to shallow models or a few spectral lines. We present a proof-of-concept study using a transformer machine learning (ML) model for multi-line, full-Stokes inversion, to infer stratified parameters from synthetic spectra produced from 3D magnetohydrodynamic simulations. We synthesise a large set of Stokes vectors using forward modelling across 15 spectral lines spanning the deep photosphere towards the chromosphere. The model maps full-Stokes input to temperature, magnetic field strength, inclination, azimuth (encoded as $\sin 2\phi$, $\cos 2\phi$), and line-of-sight velocity as a function of optical depth. The transformer incorporates an attention mechanism that allows the model to focus on the most informative regions of the spectrum for each inferred parameter, and uses positional embedding to encode wavelength and depth order. We benchmark it against a multilayer perceptron (MLP), test robustness to noise, and assess generalisation. The transformer outperforms the MLP, especially in the higher layers and for magnetic parameters, yielding higher correlations and more regularised stratifications. The model retains strong performance across a range of noise levels typical for real observations, with magnetic parameter inference degrading predictably while temperature and velocity remain stable. We establish transformer architectures as a powerful tool for spectropolarimetric regression. This approach paves the way for analysis of large datasets from large solar telescopes.
@@ 1 Introduction
Inverting the Stokes vector to retrieve physical conditions in the solar atmosphere is fundamental for understanding solar magnetism (J. C. del Toro Iniesta & B. Ruiz Cobo, 2016). While traditional inversion methods based on radiative transfer are physically grounded, they are also computationally intensive and sensitive to initialisation and noise. Crucially, the inversion problem is well suited to machine learning (ML): it is a supervised regression task with known forward physics, allowing large synthetic training sets with perfect labels, and a well-defined mapping from input (Stokes vectors as a function of wavelength) to output (atmospheric parameters as a function of geometric height or optical depth). These properties make spectropolarimetric inversion an ideal testbed for evaluating ML architectures, especially those that can handle high-dimensional inputs and uncertain mappings.
ML has been widely applied in solar physics, particularly to speed up the retrieval of atmospheric parameters from observed Stokes vectors (A. Asensio Ramos et al., 2023). A common approach involves using inversions of a small subset of Stokes vectors and associated atmospheric models as training data for a Multi-Layer Perceptron (MLP) or Convolutional Neural Network (CNN), which can then rapidly provide atmospheric models for the rest of the dataset (I. Milić & R. Gafeira, 2020). However, a major challenge is providing a diverse enough training set to achieve sufficient generalisation, necessitating large data volumes (K. E. Yang et al., 2024). R. Gafeira et al. (2021) improved the accuracy of their inversions by using models produced from CNNs as initial conditions for traditional inversions. H. Socas-Navarro & A. Asensio Ramos (2021) used only forward synthesis, an MLP, and randomly generated atmospheric models to infer target parameters, demonstrating generalisation to unseen Hinode data. Another recent advance is the use of neural networks influenced by physical equations, such as the approach taken by R. Jarolim et al. (2025) and C. U. Keller (2025).
Transformers, originally developed for natural language processing (A. Vaswani et al., 2017), have demonstrated success across many domains due to their ability to model long-range dependencies and flexible input structures. By applying a transformer-based architecture to the spectropolarimetric inversion problem, we aim to capture subtle spectral features and inter-wavelength correlations while enabling greater interpretability through attention mechanisms.
@@ 1.1 Aims and hypothesis
    Implement a transformer model that takes full Stokes vectors as input and outputs stratified atmospheric parameters (e.g., temperature, magnetic field vector, line-of-sight velocity) as a function of optical depth [^1] There is nothing in our transformer framework that prevents operation on a geometric height scale instead, as long as the forward synthesis algorithm employed operates on such a scale..
@@
The move from MLPs or CNNs to Transformers fundamentally changes what the model can learn. Transformers allow the model to attend to non-local spectral features across the ordered line profile, which is especially important in spectropolarimetry where physical information (like Zeeman splitting or velocity gradients) can manifest in subtle, non-local ways. Unlike MLPs, which treat each wavelength as a flattened feature and do not consider their order, Transformers can dynamically learn the relevant context. To our knowledge, this is the first application of an attention-based (transformer) neural network to the problem of spectropolarimetric regression in solar physics. Our central hypothesis is that a transformer-based architecture will achieve more accurate spectropolarimetric inversions than an MLP baseline, due to its ability to model spectral order (via learned positional embeddings) and non-local spectral dependencies (via attention mechanisms).
@@ 2.1 Simulation snapshots
For this study, we employ three-dimensional snapshots generated with the MANCHA code (E. Khomenko et al., 2017), as described by M. Modestov et al. (2024). By choosing a quiet Sun simulation, we deliberately select one of the most challenging regimes for spectropolarimetric regression, where polarisation signals, particularly in the chromosphere, are weak. The simulation starts from an initially field-free state, allowing the Biermann battery effect and a local turbulent dynamo to self-consistently generate magnetic fields reaching strengths comparable to those observed in the quiet Sun. These simulation snapshots have also been used in previous works such as C. Quintero Noda et al. (2023); R. J. Campbell et al. (2025). The computational domain covers an area of $11.52\times 11.52$ Mm2 (resolved with $1152\times 1152$ pixels2), and extends vertically from $1$ Mm beneath the photosphere, with a $1400$ km upper limit. The grid spacing is $10$ km in the horizontal directions and $7$ km vertically. The mean surface magnetic field strength in the saturated regime is approximately $100$ G.
To assess the generalisation performance of the ML models, we train on a given snapshot and evaluate inference on a different snapshot obtained 90 seconds later, thereby guaranteeing that the evaluation is performed on physically independent data not encountered during training.
The code used in this project to train the machine learning models is open source and is available on GitHub222www.github.com/r-j-campbell/SINN-inversions.
@@ 2.2 Forward synthesis
The simulation outputs are provided on a three-dimensional grid in geometrical height (km). To facilitate forward synthesis with the DeSIRe code, which requires an optical depth scale, we converted the atmospheric stratifications to the required log($\tau$) grid. This was done using the modelador.x utility distributed with DeSIRe, following the methodology outlined in R. J. Campbell et al. (2021).
@@
Figure 1 shows the mean intensity profile across a full snapshot. The diagnostic lines we synthesised are shown in Table 1. The Fe I $525$ nm, $630$ nm, and $1565$ nm line pairs, and the Fe I $846.84$ nm line, are synthesised in local thermodynamic equilibrium (LTE); all others are computed in non-LTE (NLTE). All spectral lines are synthesised with a spectral sampling of $20$ $\mathrm{\AA}$ per pixel. All lines were synthesised under the assumption of complete redistribution (CRD) when relevant. The spectral lines were synthesised with an observation angle, $\mu$, of $1.0$.
@@
Mg I b2 517.3 nm $-0.38616$ 2.7166 $3P_{1.0}-3S_{1.0}$ Upper photosphere/lower chromosphere
Na I D2 589 nm $+0.10106$ 0.000 $2S_{0.5}-2P_{1.5}$ Upper photosphere
Na I D1 590 nm $-0.18402$ 0.000 $2S_{0.5}-2P_{0.5}$ Upper photosphere
K I D1 770 nm $-0.16876$ 0.000 $2S_{0.5}-2P_{0.5}$ Upper photosphere
K I D2 766 nm $+0.1345$ 0.000 $2S_{0.5}-2P_{1.5}$ Upper photosphere
Ca II 849.8 nm $-1.31194$ 1.6924 $2D_{1.5}-2P_{1.5}$ Mid chromosphere (core)
Ca II 854.2 nm $-0.36199$ 1.7000 $2D_{2.5}-2P_{1.5}$ Mid chromosphere (core)
Fe I 525.02 nm $-4.938$ 0.121 $5D_{0.0}-7D_{1.0}$ Photosphere
Fe I 525.06 nm $-2.181$ 2.198 $5P_{2.0}-5P_{3.0}$ Photosphere
Fe I 630.15 nm $-0.718$ 3.654 $5P_{2.0}-5D_{2.0}$ Photosphere
Fe I 630.23 nm $-1.131$ 3.686 $5P_{1.0}-5D_{0.0}$ Photosphere
Fe I 846.84 nm $-2.072$ 2.221 $5P_{1.0}-5P_{1.0}$ Photosphere
Ti I 846.85 nm $-0.993$ 1.888 $3G_{5.0}-3F_{4.0}$ Photosphere
Fe I 1564.85 nm $-0.652$ 5.426 $7D_{1.0}-7D_{1.0}$ Deep photosphere
Fe I 1565.29 nm $+0.001$ 6.246 $7D_{5.0}-(6D_{4.5})f2k_{4.0}$ Photosphere
@@ 3.2 Transformer architecture
We implement a transformer-based neural network to perform spectropolarimetric inversions, predicting physical parameters of the solar atmosphere from observed Stokes profiles. The model ingests full Stokes vectors across multiple wavelengths and outputs stratified atmospheric parameters at all optical depth points. The goal is to replace pixel-by-pixel, node-based inversions with a model capable of learning complex, non-local spectral relationships across the Stokes vector at all wavelengths. The transformer architecture was originally developed for natural language processing (NLP) tasks, where it models sequences of words and learns the contextual meaning of each word based on its surrounding words. In that setting, attention mechanisms allow the model to focus on the most relevant parts of the sentence when generating or interpreting text. Analogously, in our application, each ”token” corresponds to the Stokes vector at a particular wavelength. The attention mechanism allows the model to learn which parts of the spectral line are most informative for inferring each physical parameter.
@@
Unlike traditional pixel-by-pixel, node-based inversion codes, our model is designed to learn complex, non-local spectral relationships that encode solar atmospheric structure. The transformer architecture is described by Fig. 2. Each input sequence corresponds to a spectral profile at a given spatial pixel, represented as a sequence of tokens. Each token corresponds to a single wavelength point and contains a 4-dimensional Stokes vector: $[I,Q,U,V]$. This forms a sequence of tokens with shape $(N_{\lambda},4)$. The model linearly projects each token into a higher-dimensional embedding space and adds learned spectral positional embeddings to preserve wavelength order. Otherwise, the model would treat the spectral sequence as unordered. [^3] The term positional encoding is often used. We adopt the term learned positional embedding here to reflect that the spectral positional information is implemented as a learned parameter matrix, rather than a fixed encoding.. During model development, we ran ablation tests with positional embedding disabled. We found that spectral positional encoding was essential, as model performance collapsed when it was removed. In contrast, depth positional encoding was less critical: the model achieved comparable validation loss without it. However, we retained it in the final model to promote depth-wise regularisation.
@@
We adopt an architectural design that enables a natural separation between the spectral domain (handled by an encoder) and the depth-dependent output domain (handled by a decoder), while preserving flexibility in the number and identity of inferred physical parameters. This encoder–decoder architecture is illustrated in Fig. 3. The embedded sequence is processed by a transformer encoder composed of stacked multi-head self-attention layers and feed-forward networks. Our encoder and decoder each consist of a tunable number of stacked layers. Each layer contains a multi-head attention mechanism followed by a position-wise feed-forward network 444“Position-wise” means that the same feed-forward network (a two-layer MLP) is applied independently to each token in the sequence. There is no interaction between sequence positions in this sublayer; all contextual mixing occurs in the attention mechanism. The feedforward sublayer (MLP) inside each transformer block uses a hidden size of 2048 (PyTorch default). (two linear layers with a dropout rate of 0.1 and ReLU activation [^5] ReLU (Rectified Linear Unit) is a common activation function defined as $\mathrm{ReLU}(x)=\max(0,x)$. Activation functions introduce non-linearity into neural networks, enabling them to approximate complex, non-linear mappings between inputs and outputs. Without them, the entire model would collapse into a linear transformation, regardless of depth.). These MLP sublayers are essential for learning non-linear combinations of spectral or depth-specific features at each token position. Each attention module uses a controllable number of heads, allowing the model to learn multiple distinct attention patterns in parallel. The outputs from each head are concatenated and linearly transformed to form the final representation for that layer. The self-attention layers compute contextualised representations of each wavelength point by comparing it to all others, enabling the model to capture subtle correlations across the spectral line profile, such as Zeeman splitting, polarisation signatures, and line asymmetries induced by velocity or magnetic gradients.
@@ 3.3 MLP baseline
To provide a simple and fast baseline, we also train a fully connected multilayer perceptron (MLP) to perform the same inversion task as the transformer. The input to the MLP is a flattened Stokes vector of shape $(4\times N_{\lambda})$, and the output is a vector of size $(6\times N_{\tau})$, representing all physical parameters at all optical depth points. As Figure 4 illustrates, the MLP architecture consists of three fully connected layers. Each hidden layer applies a linear transformation followed by a Sigmoid Linear Unit (SiLU) activation; the final output layer is linear.
This architecture processes each pixel’s Stokes vector in isolation, with no internal representation of spectral structure, positional embedding, or attention mechanism. The MLP treats the flattened input as an unordered list, lacking any explicit or implicit encoding of wavelength order, spectral continuity, or the ordering of optical depth points in the output. As a result, it cannot exploit non-local or context-dependent features in the data, and serves as a point of comparison for more expressive sequence-based models. While the MLP baseline trains significantly faster than the transformer model, it lacks the capacity to learn relationships across the spectrum, between neighbouring pixels, or across optical depths.
Figure 4: Flow diagram of the MLP inversion model. The entire input spectrum is first flattened and then passed through three fully connected layers, producing the full set of atmospheric parameters at all depths as output. The model lacks any notion of spectral order, structure, or physical relationships between wavelengths.

### v2

@@ Abstract
Solar spectropolarimetric inversion—inferring atmospheric conditions from the Stokes vector—is a key diagnostic tool for understanding solar magnetism, but traditional inversion methods are computationally expensive and sensitive to local minima. Advances in artificial intelligence (AI) offer faster solutions, but are often restricted to shallow models or a few spectral lines. We present a proof-of-concept study using a transformer machine learning (ML) model for multi-line, full-Stokes inversion, to infer stratified parameters from synthetic spectra produced from 3D magnetohydrodynamic simulations. We synthesise a large set of Stokes vectors using forward modelling across 15 spectral lines spanning the deep photosphere towards the chromosphere. The model maps full-Stokes input to temperature, magnetic field strength, inclination, azimuth (encoded as $\sin 2\phi$, $\cos 2\phi$), and line-of-sight velocity as a function of optical depth. The transformer incorporates an attention mechanism that allows the model to focus on the most informative regions of the spectrum for each inferred parameter, and uses positional embedding to encode wavelength and depth order. We benchmark it against a multilayer perceptron (MLP), test robustness to noise, and assess generalisation. The transformer outperforms the MLP, especially in the higher layers and for magnetic parameters, yielding higher correlations and more regularised stratifications. The model retains strong performance across a range of noise levels typical for real observations, with magnetic parameter inference degrading predictably while temperature and velocity remain stable. We establish transformer architectures as a powerful tool for spectropolarimetric inversion. This approach paves the way for analysis of large datasets from large solar telescopes.
@@ 1 Introduction
Inverting the Stokes vector to retrieve physical conditions in the solar atmosphere is fundamental for understanding solar magnetism (J. C. del Toro Iniesta & B. Ruiz Cobo, 2016). Inversion is the inference of depth-dependent atmospheric parameters from the wavelength dependence of the polarized spectrum in spectral lines and corresponding continua. While traditional inversion methods based on radiative transfer are physically grounded, they are also computationally intensive and sensitive to initialisation and noise (A. Asensio Ramos et al., 2012). Crucially, the inversion problem is well suited to machine learning (ML): it can be framed as a supervised regression task with known forward physics, allowing large synthetic training sets with perfect labels, and a deterministic mapping from input (Stokes vectors as a function of wavelength) to output (atmospheric parameters as a function of geometric height or optical depth). These properties make spectropolarimetric inversion an ideal application for ML architectures, especially those that can handle high-dimensional inputs and uncertain mappings.
ML has been widely applied in solar physics, particularly to speed up the retrieval of atmospheric parameters from observed Stokes vectors (A. Asensio Ramos & C. J. D´ıaz Baso, 2019; A. Asensio Ramos et al., 2023). A common approach involves using inversions of a small subset of Stokes vectors and associated atmospheric models as training data for a Multi-Layer Perceptron (MLP) or Convolutional Neural Network (CNN), which can then rapidly provide atmospheric models for the rest of the dataset (I. Milić & R. Gafeira, 2020). However, a major challenge is providing a diverse enough training set to achieve sufficient generalisation, necessitating large data volumes (K. E. Yang et al., 2024). R. Gafeira et al. (2021) improved the accuracy of their inversions by using models produced from CNNs as initial conditions for traditional inversions. H. Socas-Navarro & A. Asensio Ramos (2021) used only forward synthesis, an MLP, and randomly generated atmospheric models to infer target parameters, demonstrating generalisation to unseen Hinode data. Another recent advance is the use of neural networks influenced by physical equations, such as the approach taken by R. Jarolim et al. (2025), C. U. Keller (2025), or C. J. D´ıaz Baso et al. (2025).
Transformers, originally developed for natural language processing (A. Vaswani et al., 2017), have demonstrated success across many domains due to their ability to model long-range dependencies and flexible input structures. By applying a transformer-based architecture to the spectropolarimetric inversion problem, we aim to capture subtle spectral features and inter-wavelength correlations.
@@ 1.1 Aims and hypothesis
    Implement a transformer model that takes full Stokes vectors as input and outputs stratified atmospheric parameters (e.g., temperature, magnetic field vector, line-of-sight velocity) as a function of optical depth.
@@
The move from MLPs or CNNs to transformers fundamentally changes what the model can learn. Transformers allow the model to attend to non-local spectral features across the ordered line profile, which is especially important in spectropolarimetry where physical information (like Zeeman splitting or velocity gradients) can manifest in subtle, non-local ways. By “non-local”, we refer to correlations or informative relationships between widely separated wavelength points. Unlike MLPs, which treat each wavelength as an independent input feature and do not have awareness of their order, transformers can dynamically learn the spectral context. By “context”, we mean features whose interpretation depends on the broader spectral profiles. To our knowledge, at the time of submission, this was the first application of an attention-based (transformer) neural network to the problem of spectropolarimetric inversion in solar physics. Shortly following submission, A. Asensio Ramos & J. de la Cruz Rodriguez (2025) applied a transformer architecture to the inversion problem with a neural model used to compress the data into a lower dimensionality. Our central hypothesis is that a transformer-based architecture will achieve more accurate spectropolarimetric inversions than an MLP baseline, due to its ability to model spectral order (via learned positional embeddings) and non-local spectral dependencies (via attention mechanisms). For our study, we choose to benchmark the transformer to an MLP because without an attention mechanism the transformer architecture reduces, albeit with some caveats, to a very deep stack of MLPs (i.e. feed-forward networks). This allows us to isolate the impact of the attention mechanism and positional embeddings on inversion performance most effectively and makes a well-tuned MLP the most natural baseline model.
@@ 2.1 Simulation snapshots
For this study, we employ three-dimensional snapshots generated with the MANCHA code (E. Khomenko et al., 2017), as described by M. Modestov et al. (2024). By choosing a quiet Sun simulation, we deliberately select one of the most challenging regimes for spectropolarimetric inversion, where polarisation signals, particularly in the chromosphere, are weak. The simulation starts from an initially field-free state, allowing the Biermann battery effect and a local turbulent dynamo to self-consistently generate magnetic fields reaching strengths comparable to those observed in the quiet Sun. These simulation snapshots have also been used in previous works such as C. Quintero Noda et al. (2023); R. J. Campbell et al. (2025). The computational domain covers an area of $11.52\times 11.52$ Mm2 (resolved with $1152\times 1152$ pixels2), and extends vertically from $1$ Mm beneath the photosphere, with a $1400$ km upper limit. The grid spacing is $10$ km in the horizontal directions and $7$ km vertically. The mean surface magnetic field strength in the saturated regime is approximately $100$ G.
To assess the generalisation performance of the ML models, we train on a given snapshot and evaluate inference on a different snapshot obtained $450$ seconds later, thereby guaranteeing that the evaluation is performed on physically independent data not encountered during training.
While the chromospheric layers in the MANCHA simulation may not fully reflect the complexity of real solar conditions, they still influence the formation of spectral lines such as Ca II $854$ nm, allowing the model to learn meaningful mappings between Stokes profiles and atmospheric parameters in those layers.
@@ 2.2 Forward synthesis
The simulation outputs are provided on a three-dimensional grid in geometrical height (km). To facilitate forward synthesis with the DeSIRe code, which requires an optical depth scale, we converted the atmospheric stratifications to the required $\log(\tau)$ grid. This was done using the modelador.x utility distributed with DeSIRe, following the methodology outlined in R. J. Campbell et al. (2021). The final grid has $81$ depth points.
@@
[…]
@@
Mg I b2 517.3 nm 0.38616-0.38616 2.7166 $3P_{1.0}-3S_{1.0}$ Upper photosphere/lower chromosphere
Na I D2 589 nm +0.10106+0.10106 0.000 $2S_{0.5}-2P_{1.5}$ Upper photosphere
Na I D1 590 nm 0.18402-0.18402 0.000 $2S_{0.5}-2P_{0.5}$ Upper photosphere
K I D1 770 nm 0.16876-0.16876 0.000 $2S_{0.5}-2P_{0.5}$ Upper photosphere
K I D2 766 nm +0.1345+0.1345 0.000 $2S_{0.5}-2P_{1.5}$ Upper photosphere
Ca II 849.8 nm 1.31194-1.31194 1.6924 $2D_{1.5}-2P_{1.5}$ Mid chromosphere (core)
Ca II 854.2 nm 0.36199-0.36199 1.7000 $2D_{2.5}-2P_{1.5}$ Mid chromosphere (core)
Fe I 525.02 nm 4.938-4.938 0.121 $5D_{0.0}-7D_{1.0}$ Photosphere
Fe I 525.06 nm 2.181-2.181 2.198 $5P_{2.0}-5P_{3.0}$ Photosphere
Fe I 630.15 nm 0.718-0.718 3.654 $5P_{2.0}-5D_{2.0}$ Photosphere
Fe I 630.23 nm 1.131-1.131 3.686 $5P_{1.0}-5D_{0.0}$ Photosphere
Fe I 846.84 nm 2.072-2.072 2.221 $5P_{1.0}-5P_{1.0}$ Photosphere
Ti I 846.85 nm 0.993-0.993 1.888 $3G_{5.0}-3F_{4.0}$ Photosphere
Fe I 1564.85 nm 0.652-0.652 5.426 $7D_{1.0}-7D_{1.0}$ Deep photosphere
Fe I 1565.29 nm +0.001+0.001 6.246 $7D_{5.0}-(6D_{4.5})f2k_{4.0}$ Photosphere
@@ 3.2 Transformer architecture
We implement a transformer-based neural network to perform spectropolarimetric inversions, predicting physical parameters of the solar atmosphere from observed Stokes profiles. The model ingests full Stokes vectors across multiple wavelengths and outputs stratified atmospheric parameters at all optical depth points. The transformer architecture was originally developed for natural language processing (NLP) tasks, where it models sequences of tokens and learns the contextual meaning of each token based on its surrounding token. In that setting, attention mechanisms allow the model to focus on the most relevant parts of the sentence when generating or interpreting text. Analogously, in our application, each “token” corresponds to the Stokes vector at a particular wavelength. The attention mechanism allows the model to learn which parts of the spectral line are most informative for inferring each physical parameter.
@@
The transformer architecture is described by Fig. 2. Each input sequence corresponds to a spectral profile at a given spatial pixel, represented as a sequence of tokens. Each token corresponds to a single wavelength point and contains a 4-dimensional Stokes vector: $[I,Q,U,V]$. This forms a sequence of tokens with shape $(N_{\lambda},4)$. The model linearly projects each token into a higher-dimensional embedding space and adds learned spectral positional embeddings to preserve wavelength order. Otherwise, the model would treat the spectral sequence as unordered. [^1] The term positional encoding is often used. We adopt the term learned positional embedding here to reflect that the spectral positional information is implemented as a learned parameter matrix, rather than a fixed encoding.. During model development, we ran ablation tests with positional embedding disabled. We found that spectral positional embedding was essential, as model performance collapsed when it was removed. In contrast, depth positional embedding was less critical: the model achieved comparable validation loss without it. However, we retained it in the final model to promote depth-wise regularisation.
@@
We adopt an architectural design that enables a natural separation between the spectral domain (handled by an encoder) and the depth-dependent output domain (handled by a decoder), while preserving flexibility in the number and identity of inferred physical parameters. This encoder–decoder architecture is illustrated in Fig. 3. The embedded sequence is processed by a transformer encoder composed of stacked multi-head self-attention layers and feed-forward networks. Our encoder and decoder each consist of a tunable number of stacked layers. Each layer contains a multi-head attention mechanism followed by a position-wise feed-forward network 222“Position-wise” means that the same feed-forward network (a two-layer MLP) is applied independently to each token in the sequence. There is no interaction between sequence positions in this sublayer; all contextual mixing occurs in the attention mechanism. The feedforward sublayer (MLP) inside each transformer block uses a hidden size of $2048$ (PyTorch default). (two linear layers with a dropout rate of 0.1 and ReLU activation [^3] ReLU (Rectified Linear Unit) is a common activation function defined as $\mathrm{ReLU}(x)=\max(0,x)$. Activation functions introduce non-linearity into neural networks, enabling them to approximate complex, non-linear mappings between inputs and outputs. Without them, the entire model would collapse into a linear transformation, regardless of depth.). These MLP sublayers are essential for learning non-linear combinations of spectral or depth-specific features at each token position. Each attention module uses a controllable number of heads, allowing the model to learn multiple distinct attention patterns in parallel. The outputs from each head are concatenated and linearly transformed to form the final representation for that layer. The self-attention layers compute contextualised representations of each wavelength point by comparing it to all others, enabling the model to capture subtle relationships across the spectral line profile.
@@ 3.3 MLP baseline
To provide a simple and fast baseline, we also train a fully connected multilayer perceptron (MLP) to perform the same inversion task as the transformer. The input to the MLP is a flattened Stokes vector of shape $(4\times N_{\lambda})$, and the output is a vector of size $(6\times N_{\tau})$, representing all physical parameters at all optical depth points. As Figure 4 illustrates, a 2-hidden-layer MLP architecture consists of three fully connected layers. Each hidden layer applies a linear transformation followed by a Sigmoid Linear Unit (SiLU) activation; the final output layer is linear. We use MLP architectures with more than two hidden layers, but Fig. 4 shows an example with only two for clarity.
This architecture processes each pixel’s Stokes vector in isolation, with no internal representation of spectral structure, positional embedding, or attention mechanism. The MLP treats the flattened input as an unordered list, lacking any explicit or implicit encoding of wavelength order, spectral continuity, or the ordering of optical depth points in the output. As a result, it cannot exploit non-local or context-dependent features in the data, and serves as a point of comparison for more expressive sequence-based models. While the MLP baseline trains significantly faster than the transformer model, it has no explicit mechanism to model relationships along the spectrum or across optical depths, relying instead on learning such dependencies implicitly from the data.
Figure 4: Flow diagram of a 2-hidden-layer MLP inversion model. The entire input spectrum is first flattened and then passed through three fully connected layers, producing the full set of atmospheric parameters at all depths as output.